OpenStudy (paige):

Differentiate x(4x+12)^1/2 to give it in the form of k(x+2)/(4x+12)^1/2.

OpenStudy (anonymous):

this is easy

OpenStudy (paige):

Help me please.

OpenStudy (anonymous):

Do you know how to take derivatives? f'(x)/g'(x) = (f'(x)g(x)-f(g)g'(x))/g'(x)^2

OpenStudy (anonymous):

that is wrong. \[g'(x)=(4x+12)^{-\frac{1}{2}}\times 4\]

OpenStudy (paige):

WHAT? O.o :S

OpenStudy (anonymous):

\[(fg)'=f'g+g'f\]

OpenStudy (anonymous):

\[f(x)=x\] \[f''(x)=1\] \[g(x)=\sqrt{4x+12}\] \[g'(x)=\frac{4}{2\sqrt{4x+12}}=\frac{2}{\sqrt{4x+12}}\]\]

OpenStudy (anonymous):

put it together in the product rule

OpenStudy (anonymous):

site is real slow. any steps you need written out?

OpenStudy (paige):

It's too complicated to read :S